OpenStudy (anonymous):

Write the values of the trigonometric functions that are defined for theta=x. csc pi a. -1 b. is undefined c. 0 d. 1

OpenStudy (ash2326):

\[ cosec\ x= \frac{1}{\sin x}\] @mathsux4real Do you know the value of \(\sin \pi\)?

OpenStudy (ash2326):

\(\sin \pi=0\) so what do you think is the answer?

OpenStudy (anonymous):

Thanks ash!